A gas has a volume of 3.0 L at 330K. What will its

volume be at 373K?


1
Expert's answer
2021-05-11T05:57:04-0400

V1 = 3.0L. T1= 330K

V2 = ? T2= 373K

V2/T2 = V1/ T1

V2/373 = 3.0/330

V2 = 3.4L